# NeuroSpex: Neuro-Guided Speaker Extraction with Cross-Modal Fusion

### v1

@@ 1 Introduction
Studies in neuroscience have revealed a strong correlation between the attended speech and the neural response it elicits [4], paving the way for auditory attention detection (AAD) from neural activities [5, 6]. Such neural responses include Electrocorticography (ECoG) [4], Magnetoencephalography (MEG) [6], and EEG [5]. Among them, EEG provides a non-invasive, comparably easy-to-wear, and affordable choice for SAA tasks, especially for neuro-steered hearing aids [7].
@@ 4.1 Dataset
EEG recordings were referenced to the average of all electrodes, band-pass filtered between 1 and 32 Hz, and downsampled to 128 Hz. This frequency range was selected to align with auditory attention detection protocols. Data normalization was performed trial-by-trial to standardize mean and variance across the dataset. In total, the dataset comprises 128 trials amounting to 12.8 hours of speech-EEG parallel data.
@@ 4.4 Model Training
All implementations utilize the PyTorch framework with distributed training and data distributed sampler techniques across 2 Nvidia RTX A6000 GPUs. To ensure reproducibility, we used random seeds to generate consistent results. The models are trained end-to-end using the Adam optimizer with an initial learning rate of 0.0001. A learning rate scheduler with a decay factor of 0.5 when the best validation loss does not improve within 5 consecutive epochs and early stopping is applied when the best validation loss does not improve within the last 25 epochs. Training is conducted for around 100 epochs or until stable training with batch size of 16.
@@
Table 2: Effect of the number of AdC blocks on the proposed model performance to find best configuration using multiple metrics evaluated on the test set.
@@ 5.3 Comparison with Baseline Models
We compare our best model (with 6 AdC blocks) with multiple baseline models as shown in Table 3 including the main architectural difference between models and evaluation metrics. The violin plots in Fig. 4 summarize the SI-SDRi for all trials in the test set for all models. The BSS (PIT) shows the best results with the least variance as it performs direct speech separation and sets an upper bound for the comparison. Our proposed model outperforms all baselines significantly (p $<$ 0.001) except BSS (PIT) on all evaluation metrics showing that it performs the speaker extraction with better signal quality perpetual quality, and intelligibility. Furthermore, the proposed NeuroSpex model contains fewer parameters compared to the ConvTasNet based NeuroHeed model but more than DPRNN counterparts due to the TCN based speaker extractor. According to violin plots, the proposed model has a better-centered distribution around a higher median and less variance compared to both NeuroHeed and BASEN baselines.

### v2

@@ 1 Introduction
Studies in neuroscience have revealed a strong correlation between the attended speech and the neural response it elicits [4], paving the way for auditory attention detection (AAD) from neural activities. Such neural responses include Electrocorticography (ECoG) [4], Magnetoencephalography (MEG) [5], and EEG [6]. Among them, EEG provides a non-invasive, comparably easy-to-wear, and affordable choice for SAA tasks, especially for neuro-steered hearing aids [7].
@@ 4.1 Dataset
EEG recordings were referenced to the average of all electrodes, band-pass filtered between 1 and 32 Hz, and downsampled to 128 Hz. This frequency range was selected to align with AAD protocols. Data normalization was performed trial-by-trial to standardize mean and variance across the dataset. In total, the dataset comprises 128 trials amounting to 12.8 hours of speech-EEG parallel data.
@@ 4.4 Model Training
All implementations utilize the PyTorch framework with distributed training and data distributed sampler techniques across 2 Nvidia RTX A6000 GPUs. To ensure reproducibility, we used random seeds to generate consistent results. The models are trained end-to-end using the Adam optimizer with an initial learning rate of 0.0001. A learning rate scheduler with a decay factor of 0.5 when the best validation loss does not improve within 5 consecutive epochs and early stopping is applied when the best validation loss does not improve within the last 25 epochs. Training is conducted for around 100 epochs or until stable training with a batch size of 16.
@@
Table 2: Impact of the number of AdC blocks on the performance of our proposed NeuroSpex, evaluated using multiple metrics on the test set
@@ 5.3 Comparison with Baseline Models
We compare our best model (with 6 AdC blocks) with multiple baseline models as shown in Table 3 including the main architectural difference between models and evaluation metrics. The violin plots in Fig. 4 summarize the SI-SDRi for all trials in the test set for all models. The BSS (PIT) shows the best results with the least variance as it performs direct speech separation and sets an upper bound for the comparison. Our proposed model outperforms all baselines significantly (p $<$ 0.001) except BSS (PIT) on all evaluation metrics showing that it performs the speaker extraction with better signal quality perpetual quality, and intelligibility. Furthermore, the proposed NeuroSpex model contains fewer parameters compared to the ConvTasNet-based NeuroHeed model but more than DPRNN counterparts due to the TCN-based speaker extractor. According to violin plots, the proposed model has a better-centered distribution around a higher median and less variance compared to both NeuroHeed and BASEN baselines.